Use cylindrical shells to find the volume of the solid that results when the region enclosed by x = y^2 and x = y is revolved about the line y= -1
1
Expert's answer
2020-09-06T18:22:15-0400


We should rotate the region above about the line y=-1.This volume can be represented as the difference between the volume V1 obtained by rotating the upper curve and the volume V2 obtained by rotating the lower curve. Two curves intersect when "y^2=y" in points "(0;0), (1;1)" .

Every volume can be represented as he sum of volumes of small cylinders with height dx and width f(x)-(-1), so the volume of such an cylinder is "\pi (f(x)+1)^2dx" Therefore, to obtain V1 and V2 we should sum the volumes of cylinders. If we consider smaller and smaller dx, the sum will turn into integral:

"V_1 = \int\limits_0^1 \pi (\sqrt{x}+1)^2dx =\pi \int\limits_0^1 (x+2\sqrt{x}+1)dx= \pi \left(\dfrac{x^2}{2} + \dfrac{4}{3}x^{3/2}+x \right) \Big|^1_0 = \dfrac{17}{6}\pi."

"V_2 = \int\limits_0^1 \pi (x+1)^2dx = \pi\left(x^3/3+x^2+x\right)\Big|^1_0 = \dfrac{7\pi}{3}."

The difference between volumes is "\dfrac{\pi}{2}."